OpenStudy (amonoconnor):

I need help in evaluating the integral of the following: [(cos(x))/((2 + sin(x))^2)]dx, from x=(-pi/2) to (pi/2). I evaluated using u-substitution, where u = (2 + sin(x)) and got a final answer of: 0. WolframAlpha got: 2/3 I have reworked the problem a few times, and cannot determine what I'm doing wrong. Can someone do this problem, verify which answer (if either of them) is correct, and if I am wrong, work through the problem with me? Any and all help is greatly appreciated! :)

OpenStudy (p0sitr0n):

so its \[\int\limits_{-\Pi/2}^{\Pi/2}\frac{ \cos x }{ 2+\sin x^{2} }\], right?

OpenStudy (amonoconnor):

Very close! It may not matter, but the bottom looks like this: (2 + sin(x))^2. (Squared is separated by parentheses)

OpenStudy (p0sitr0n):

oh kk sorry

OpenStudy (anonymous):

\[\int\limits_{\frac{- \pi }{ 2 }}^{\frac{ \pi }{ 2 }}\left( 2+\sin x \right)^{-2}\cos x~dx\] \[=\frac{ -1 }{ 2+\sin x },from ~\frac{ -\pi }{ 2 }\rightarrow \frac{ \pi }{ 2 }\] now you can solve. i am leaving now.

OpenStudy (p0sitr0n):

so its I=- 1/(2+sinx) from -PI/2, PI/2 -1/(2+1)-(-1/2-1)=-1/3+1=2/3

OpenStudy (p0sitr0n):

just careful with the signs

OpenStudy (amonoconnor):

Oh, I see my mistake :D *eraser flying....

OpenStudy (p0sitr0n):

xD

OpenStudy (amonoconnor):

...I'm almost done reworking:)

OpenStudy (p0sitr0n):

ah kk nice

OpenStudy (amonoconnor):

Man do I feel stupid now.... :) I got it, +2/3. Thank you very much!